Premise:
Television sets. What does it mean to refer to a 20-in TV set or a 25-in TV set? Such units refer to the diagonal of the screen. A 10-in TV set also has a width of 8 inches.
Question:
What is the height?
Please explain how you arrived at this answer. Thanks!

Answer by stanbon(75887)

What does it mean to refer to a 20-in TV set or a 25-in TV set?
Such units refer to the diagonal of the screen.
A 10-in TV set also has a width of 8 inches.
Question:
What is the height?
Draw a rectangle that has a width = 8" and a diagonal = 10"
-----
Let the height be "x".
Use Pythagoras to find "x":
Equation:
10^2 = x^2 + 8^2
100 = x^2 + 64
x^2 = 36
x = 6"
Ans: height = 6"
=======================
Cheers,
Stan H.